OpenStudy (anonymous):

Convert the equation x – y = -3 to slope-intercept form. y = x – 3 y = -x + 3 y = -x – 3 y = x + 3

OpenStudy (anonymous):

rearrange the eqn x - y = -3 in the form y = mx + c got it ??

OpenStudy (anonymous):

what will mx and c be?

OpenStudy (anonymous):

you got nothing to do with mx and c x- y = -3 -y = -3 - x multiply by - sign you will get ur answer

OpenStudy (anonymous):

y=3+x?

OpenStudy (anonymous):

yaaa..

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

welcome